OpenStudy (anonymous):

If 13.8 moles of carbon (C), with a density of 2.26 g/mL, are needed for a chemical reaction, how many milliliters of carbon are needed? 73.3 mL C 16.5 mL C 6.11 mL C 2.60 mL C (I know this isn't math but the chem section is dead and I'm desperate for some help) @amistre64 @ganeshie8 @abhisar can you guys help?

OpenStudy (abhisar):

Molar mass of carbon is 12, so mass of 13.8 mole carbon = 13.8*12=165 g. Now density is 2.26g/ml so volume of carbon required = 165/2.26 = ?

OpenStudy (anonymous):

73.01

OpenStudy (anonymous):

or 73.008

OpenStudy (abhisar):

\(\huge\checkmark\)

OpenStudy (anonymous):

ohh so that's the answeR?

OpenStudy (anonymous):

Its A?

OpenStudy (abhisar):

SO the correct choice will be ?

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @brriiarr Its A? \(\huge\checkmark\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

thanks so much<3

OpenStudy (abhisar):

\(\rlap{\color{blue}{\huge\bigstar}\huge \color{blue}{ \text{You're Most Welcome! }}\color{blue}\bigstar}{\; \color{aqua}{\huge\bigstar}\huge \color{aqua}{\text{You're Most Welcome! }}\color{aqua}\bigstar}\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)